OpenStudy (anonymous):

Find three consecutive odd numbers whose sum is 1767.

OpenStudy (anonymous):

use: 2n + 1, 2n + 3, 2n + 5

OpenStudy (anonymous):

the answers are 881,882,883

OpenStudy (anonymous):

882 isnt odd

OpenStudy (anonymous):

that what i got when i use the equation 2n+3

OpenStudy (anonymous):

show me what u did?

OpenStudy (anonymous):

2n+3=1767 -3 -3 2n=1767 n=882

OpenStudy (anonymous):

\[[2n+1]+[2n+3]+[2n+5]=1767\] \[6n+9=1767\] \[n=293\] Now plug that n value back into your formula to find the first value

hartnn (hartnn):

The 'sum' is 1767. so 2n+1+2n+3+2n+5=1767 so 6n+9=1767 can u find n from here?

OpenStudy (anonymous):

yes